### v1

# The ratio monotonicity of Eulerian-type polynomials

Abstract
This paper is motivated by determining the location of modes of some unimodal Eulerian-type polynomials. The notion of ratio monotonicity was introduced by Chen-Xia when they investigated the $q$-derangement numbers. Let $(f_{n}(x))_{n\geqslant 0}$ be a sequence of real polynomials satisfying the Eulerian-type recurrence relation
$f_{n+1}(x)=(anx+bx+c)f_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}f_{n}(x),~f_{0}(x)=1,$
where $a,b$ and $c$ are nonnegative integers. Assume that $\deg f_{n}(x)=n$. Setting $g_{n}(x)=x^{n}f_{n}\left(\frac{1}{x}\right)$, we have
$g_{n+1}(x)=(anx+b+cx)g_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}g_{n}(x),$
We find that if $a+c\geqslant b\geqslant c>0$, then $f_{n}(x)$ is bi-gamma-positive and $g_{n}(x)$ is ratio monotone. As applications, we discover the ratio monotonicity of several Eulerian-type polynomials, including the $({\rm exc\,},{\rm cyc\,})$ $q$-Eulerian polynomials, the $1/k$-Eulerian polynomials, a kind of generalized Eulerian polynomials studied by Carlitz-Scoville, the $(\operatorname{des}_{B},{\rm neg\,})$ $q$-Eulerian polynomials over the hyperoctahedral group and the $r$-colored Eulerian polynomials. In particular, let $A_{n}(x,q)$ be the $({\rm exc\,},{\rm cyc\,})$ $q$-Eulerian polynomials, we find that the polynomials $x^{n-1}A_{n}(1/x,q)$ are ratio monotone when $0<q\leqslant 1$, while $A_{n}(x,q)$ are ratio monotone when $1\leqslant q\leqslant 2$.
keywords:
Ratio monotonicity , Eulerian polynomials , Recurrence systems, Unimodality
MSC:
[2010] 05A05, 05A15
1 Introduction
The types $A$ and $B$ Eulerian polynomials can be respectively defined by
$\begin{split}A_{n+1}(x)&=(nx+1)A_{n}(x)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}A_{n}(x),\\ B_{n+1}(x)&=(2nx+x+1)B_{n}(x)+2x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}B_{n}(x),\end{split}$
with $A_{0}(x)=B_{0}(x)=1$ (see [7, 13]). There has been much recent work on Eulerian-type recursions, see [3, 17, 22, 31] for instance. For example, Hwang-Chern-Duh [17] considered the general recurrence relation:
$\mathcal{P}_{n}(x)=(\alpha(x)n+\gamma(x))\mathcal{P}_{n-1}(x)+\beta(x)(1-x)\frac{\mathrm{d}}{\mathrm{d}x}\mathcal{P}_{n-1}(x),$
where $\mathcal{P}_{0}(x),\alpha(x),\beta(x)$ and $\gamma(x)$ are given functions of $x$. They studied the limiting distribution of the coefficients of $\mathcal{P}_{n}(x)$. Very recently, Liu-Yan [19] considered interlacing properties related to the polynomials $f_{n}(x)$ satisfying the recurrence relation
$f_{n}(x)=(a_{1}n+a_{2}+(b_{1}n+b_{2})x)f_{n-1}(x)+c_{1}(n-1)xf_{n-2}(x)+x(a_{3}+b_{3}x)\frac{\mathrm{d}}{\mathrm{d}x}f_{n-1}(x),$
with $f_{0}(x)=1$. Unimodal and log-concave polynomials arise often in combinatorics and other branches in mathematics, but to determine the location of modes may be a difficult challenge, see [4, 6] for surveys on this topic. This paper is motivated by determining the location of modes of some unimodal Eulerian-type polynomials.
Let $a(x)=\sum_{i=0}^{n}a_{i}x^{i}$ be a polynomial with nonnegative coefficients. We say that $a(x)$ is unimodal if $a_{0}\leqslant a_{1}\leqslant\cdots\leqslant a_{k-1}\leqslant a_{k}\geqslant a_{k+1}\geqslant\cdots\geqslant a_{n}$ for some $k$, where the index $k$ is called the mode of $a(x)$. It is log-concave if $a_{i}^{2}\geqslant a_{i-1}a_{i+1}$ for all $1\leqslant i\leqslant n-1$. Clearly, a log-concave sequence with no internal zeros is unimodal, but the converse is not true. The polynomial $a(x)$ is said to be gamma-positive if there exist nonnegative numbers $\gamma_{k}$ such that
$a(x)=\sum_{k=0}^{\lfloor{n}/{2}\rfloor}\gamma_{k}x^{k}(1+x)^{n-2k}.$
Clearly, gamma-positivity is a property that polynomials with symmetric coefficients may have, which implies their unimodality and symmetry, see [1] for a survey on this subject.
We say that $a(x)$ is spiral if $a_{n}\leqslant a_{0}\leqslant a_{n-1}\leqslant a_{1}\leqslant\cdots\leqslant a_{\lfloor n/2\rfloor}$. Following [26, Definition 2.9], the polynomial $a(x)$ is alternatingly increasing if $a_{0}\leqslant a_{n}\leqslant a_{1}\leqslant a_{n-1}\leqslant\cdots\leqslant a_{\lfloor{(n+1)}/{2}\rfloor}$. Clearly, if $a(x)$ is spiral and $\deg a(x)=n$, then $x^{n}a(1/x)$ is alternatingly increasing, and vice versa. The notion of ratio monotonicity for polynomials was introduced by Chen-Xia [12]. We say that $a(x)$ is ratio monotone if
$\frac{a_{n}}{a_{0}}\leqslant\frac{a_{n-1}}{a_{1}}\leqslant\cdots\leqslant\frac{a_{n-i}}{a_{i}}\leqslant\frac{a_{n-i-1}}{a_{i+1}}\leqslant\cdots\leqslant\frac{a_{n-\left[\frac{n-1}{2}\right]}}{a_{\left[\frac{n-1}{2}\right]}}\leqslant 1$ (1)
and
$\frac{a_{0}}{a_{n-1}}\leqslant\frac{a_{1}}{a_{n-2}}\leqslant\cdots\leqslant\frac{a_{i-1}}{a_{n-i}}\leqslant\frac{a_{i}}{a_{n-i-1}}\leqslant\cdots\leqslant\frac{a_{\left[\frac{n}{2}\right]-1}}{a_{n-\left[\frac{n}{2}\right]}}\leqslant 1.$ (2)
Ratio monotonicity implies the log-concavity and spiral property. From (1) and (2), we have
$\frac{a_{i+1}}{a_{i}}\leqslant\frac{a_{n-i-1}}{a_{n-i}}\leqslant\frac{a_{i}}{a_{i-1}}.$
In [10, 12], Chen-Xia gave elegant proofs of the ratio monotonicity of Boros-Moll polynomials and $q$-derangement numbers. Let $P(x)$ be a polynomial with nonnegative and nondecreasing coefficients. Chen-Yang-Zhou [11] proved the ratio monotone property of $P(x+1)$. Since then, there has been much attention on the ratio monotone property of enumerative polynomials. For example, Su-Sun [29] established ratio monotonicity of the coordinator polynomials of the root lattice of type $B_{n}$.
Let $f(x)$ be a polynomial of degree $n$. There is a unique symmetric decomposition $f(x)=a(x)+xb(x)$, where
$a(x)=\frac{f(x)-x^{n+1}f(1/x)}{1-x},~b(x)=\frac{x^{n}f(1/x)-f(x)}{1-x}.$ (3)
Clearly, $a(x)$ and $b(x)$ are both symmetric. Moreover, $\deg a(x)=1+\deg b(x)$ or $b(x)=0$. Following Brändén-Solus [5], we call the ordered pair of polynomials $(a(x),b(x))$ the symmetric decomposition of $f(x)$. According to [21, Definition 1.2], if $a(x)$ and $b(x)$ are both gamma-positive, then $f(x)$ is said to be bi-gamma-positive. Bi-gamma-positivity implies alternatingly increasing property and gamma-positivity can be seen as a degeneration of bi-gamma-positivity. For example, the polynomial $1+10x+4x^{2}$ is bi-gamma-positive, since
$1+10x+4x^{2}=(1+7x+x^{2})+x(3+3x)=((1+x)^{2}+5x)+x(3(1+x)).$
We can now present the main result of this paper.
Theorem 1.
Let $(f_{n}(x))_{n\geqslant 0}$ be a sequence of real polynomials satisfying the Eulerian-type recurrence
$f_{n+1}(x)=(anx+bx+c)f_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}f_{n}(x),$ (4)
with $f_{0}(x)=1$, where $a,b$ and $c$ are nonnegative integers. Suppose that $\deg f_{n}(x)=n$. If $a+c\geqslant b\geqslant c>0$, then $f_{n}(x)$ is bi-gamma-positive and $x^{n}f_{n}(1/x)$ is ratio monotone.
2 Proof of Theorem 1
2.1 A proof of the bi-gamma-positivity of $f_{n}(x)$
Applying the formula (3), we have $f_{n}(x)=a_{n}(x)+xb_{n}(x)$, where
$a_{n}(x)=\frac{f_{n}(x)-x^{n+1}f_{n}(1/x)}{1-x},~b_{n}(x)=\frac{x^{n}f_{n}(1/x)-f_{n}(x)}{1-x}.$
Note that $x^{n}f_{n}(1/x)=x^{n}a_{n}(1/x)+x^{n-1}b_{n}(1/x)=a_{n}(x)+b_{n}(x)$. Setting $g_{n}(x)=x^{n}f_{n}(1/x)$, it is easy to verify that
$g_{n+1}(x)=(anx+b+cx)g_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}g_{n}(x).$ (5)
Using (4) and (5), $f_{n}(x)=a_{n}(x)+xb_{n}(x)$ and $g_{n}(x)=a_{n}(x)+b_{n}(x)$ can be rewritten as
$a_{n+1}(x)+xb_{n+1}(x)$ $=(anx+bx+c)a_{n}(x)+x(a(n-1)x+bx+a+c)b_{n}(x)+$
$ax(1-x)\left(\frac{\mathrm{d}}{\mathrm{d}x}a_{n}(x)+x\frac{\mathrm{d}}{\mathrm{d}x}b_{n}(x)\right);$
$a_{n+1}(x)+b_{n+1}(x)$ $=(anx+b+cx)a_{n}(x)+(anx+b+cx)b_{n}(x)+$
$ax(1-x)\left(\frac{\mathrm{d}}{\mathrm{d}x}a_{n}(x)+\frac{\mathrm{d}}{\mathrm{d}x}b_{n}(x)\right).$
In view of
$(1-x)a_{n+1}(x)=a_{n+1}(x)+xb_{n+1}(x)-x\left(a_{n+1}(x)+b_{n+1}(x)\right),$
$(x-1)b_{n+1}(x)=a_{n+1}(x)+xb_{n+1}(x)-\left(a_{n+1}(x)+b_{n+1}(x)\right),$
we obtain the following recurrence system
$\left\{\begin{array}[]{ll}a_{n+1}(x)&=(anx+c+cx)a_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}a_{n}(x)+(a-b+c)xb_{n}(x),\\ b_{n+1}(x)&=(anx-ax+bx+b)b_{n}(x)+ax(1-x)\frac{\mathrm{d}}{\mathrm{d}x}b_{n}(x)+(b-c)a_{n}(x),\end{array}\right$. (6)
with the initial conditions $a_{0}(x)=1$ and $b_{0}(x)=0$. In particular, we have
$a_{1}(x)=c+cx,~b_{1}(x)=b-c;$
$a_{2}(x)=c^{2}+(ab-b^{2}+ac+2bc+c^{2})x+c^{2}x^{2},~b_{2}(x)=(b^{2}-c^{2})(1+x).$
Assume that
$a_{n}(x)=\sum_{k\geqslant 0}\alpha_{n,k}x^{k}(1+x)^{n-2k},~b_{n}(x)=\sum_{k\geqslant 0}\beta_{n,k}x^{k}(1+x)^{n-1-2k}.$
Using (6), it is routine to verify that
$\left\{\begin{array}[]{ll}\alpha_{n+1,k}=(ak+c)\alpha_{n,k}+2a(n-2k+2)\alpha_{n,k-1}+(a-b+c)\beta_{n,k-1},\\ \beta_{n+1,k}=(ak+b)\beta_{n,k}+2a(n-2k+1)\beta_{n,k-1}+(b-c)\alpha_{n,k},\end{array}\right$. (7)
with $\alpha_{0,0}=1,~\beta_{0,0}=0$ and $\alpha_{0,k}=\beta_{0,k}=0$ for $k\geqslant 1$. Therefore, when $a+c\geqslant b\geqslant c>0$, we see that $f_{n}(x)$ is bi-gamma-positive, and when $a>0$ and $b=c$, $f_{n}(x)$ is gamma-positive. ∎
Suppose that $p(x)$ and $q(x)$ have only real zeros, the zeros of $p(x)$ are $\xi_{1}\leqslant\cdots\leqslant\xi_{n}$, and that those of $q(x)$ are $\theta_{1}\leqslant\cdots\leqslant\theta_{m}$. Following [14], we say that $p(x)$ interlaces $q(x)$ if $\deg q(x)=1+\deg p(x)$ and the zeros of $p(x)$ and $q(x)$ satisfy $\theta_{1}\leqslant\xi_{1}\leqslant\theta_{2}\leqslant\cdots\leqslant\xi_{n}\leqslant\theta_{n+1}$. We use the notation $p(x)\prec q(x)$ for $p(x)$ interlaces $q(x)$. Combining (6) and [18, Theorem 2.3], one can immediately get the following result.
Proposition 2.
Let $\{f_{n}(x)\}_{n\geqslant 0}$ be a sequence of polynomials satisfying (4). Then $f_{n}(x)$ has only real zeros. Suppose that $(a_{n}(x),b_{n}(x))$ is the symmetric decomposition of $f_{n}(x)$. Then both $a_{n}(x)$ and $b_{n}(x)$ have only real zeros and $b_{n}(x)\prec a_{n}(x)$.
2.2 A proof of the ratio monotonicity of $f_{n}(x)$
To prove the ratio monotonicity of $f_{n}(x)$, we need the following lemma.
Lemma 3.
Suppose that $a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}$ are positive numbers satisfying
$\frac{a_{1}}{a_{2}}\leqslant\frac{a_{3}}{a_{4}}\leqslant\frac{a_{5}}{a_{6}},~\frac{a_{2}}{a_{3}}\leqslant\frac{a_{4}}{a_{5}},~a_{3}\leqslant a_{5},~a_{4}\leqslant a_{6},$ (8)
then we have
$\frac{\lambda_{1}a_{1}+(\lambda-\lambda_{1})a_{3}}{\lambda_{2}a_{2}+(\lambda-\lambda_{2})a_{4}}\leqslant\frac{\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}+\mu(a_{5}-a_{3})}{\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}+\mu(a_{6}-a_{4})},$
where $0<\lambda_{2}\leqslant\lambda_{1}\leqslant\lambda$ and $\mu\geqslant 0$.
Proof.
Note that $a_{1}(a_{4}+a_{6})-a_{2}(a_{3}+a_{5})=(a_{1}a_{4}-a_{2}a_{3})+(a_{1}a_{6}-a_{2}a_{5})\leqslant 0$. We get
$\frac{a_{1}}{a_{2}}\leqslant\frac{a_{3}+a_{5}}{a_{4}+a_{6}}.$ (9)
We now show that
$\frac{\lambda_{1}a_{1}+(\lambda-\lambda_{1})a_{3}}{\lambda_{2}a_{2}+(\lambda-\lambda_{2})a_{4}}\leqslant\frac{\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}}{\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}}.$ (10)
It follows from (8) that
$a_{1}a_{4}\leqslant a_{2}a_{3},~a_{3}a_{6}\leqslant a_{4}a_{5},~a_{1}a_{6}\leqslant a_{2}a_{5},~\frac{a_{1}}{a_{4}}=\frac{a_{1}}{a_{2}}\frac{a_{2}}{a_{4}}\leqslant\frac{a_{5}}{a_{6}}\frac{a_{3}}{a_{5}}=\frac{a_{3}}{a_{6}}.$
Therefore, we obtain
$\left(\lambda_{1}a_{1}+(\lambda-\lambda_{1})a_{3}\right)\left(\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}\right)-\left(\lambda_{2}a_{2}+(\lambda-\lambda_{2})a_{4}\right)\left(\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}\right)$
$=\lambda_{1}\lambda_{2}(a_{1}a_{4}-a_{2}a_{3})+(\lambda-\lambda_{1})(\lambda-\lambda_{2})(a_{3}a_{6}-a_{4}a_{5})+$
$\lambda_{1}(\lambda-\lambda_{2})a_{1}a_{6}-\lambda(\lambda_{1}-\lambda_{2})a_{3}a_{4}-\lambda_{2}(\lambda-\lambda_{1})a_{2}a_{5}$
$=\lambda_{1}\lambda_{2}(a_{1}a_{4}-a_{2}a_{3})+(\lambda-\lambda_{1})(\lambda-\lambda_{2})(a_{3}a_{6}-a_{4}a_{5})+$
$(\lambda_{1}\lambda-\lambda_{2}\lambda+\lambda_{2}\lambda-\lambda_{1}\lambda_{2})a_{1}a_{6}-\lambda(\lambda_{1}-\lambda_{2})a_{3}a_{4}-\lambda_{2}(\lambda-\lambda_{1})a_{2}a_{5}$
$=\lambda_{1}\lambda_{2}(a_{1}a_{4}-a_{2}a_{3})+(\lambda-\lambda_{1})(\lambda-\lambda_{2})(a_{3}a_{6}-a_{4}a_{5})+$
$\lambda(\lambda_{1}-\lambda_{2})(a_{1}a_{6}-a_{3}a_{4})+\lambda_{2}(\lambda-\lambda_{1})(a_{1}a_{6}-a_{2}a_{5})\leqslant 0,$
which yields (10), as desired.
Note that
$\left(\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}\right)a_{6}-\left(\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}\right)a_{5}$
$=\lambda_{1}a_{3}a_{6}-\lambda_{2}a_{4}a_{5}+(\lambda_{2}-\lambda_{1})a_{5}a_{6}$
$=\lambda_{1}a_{6}(a_{3}-a_{5})+\lambda_{2}a_{5}(a_{6}-a_{4})$
$\leqslant\lambda_{1}a_{6}(a_{3}-a_{5})+\lambda_{1}a_{5}(a_{6}-a_{4})$
$=\lambda_{1}(a_{3}a_{6}-a_{4}a_{5})\leqslant 0.$
So we obtain
$\frac{\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}}{\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}}\leqslant\frac{a_{5}}{a_{6}}\leqslant\frac{a_{5}-a_{3}}{a_{6}-a_{4}},$ (11)
where the last inequality can be easily verified. In conclusion, we get
$\frac{\lambda_{1}a_{1}+(\lambda-\lambda_{1})a_{3}}{\lambda_{2}a_{2}+(\lambda-\lambda_{2})a_{4}}\leqslant\frac{\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}}{\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}}\leqslant\frac{a_{5}-a_{3}}{a_{6}-a_{4}}.$
Comparing with (8) and (9), we immediately get the desired result. This completes the proof. ∎
A proof of the ratio monotonicity of $f_{n}(x)$:
Proof.
Comparing (4) and (5), we see that the following two statements are equivalent:
  • $(i)$
    If $a+c\geqslant b\geqslant c>0$, then $x^{n}f_{n}(1/x)$ is ratio monotone;
  • $(ii)$
    If $a+b\geqslant c\geqslant b>0$, then $f_{n}(x)$ is ratio monotone.
Let $f_{n}(x)=\sum_{i=0}^{n}f_{n,i}x^{i}$. It follows from (4) that
$f_{n+1,i}=(ai+c)f_{n,i}+(a(n-i)+a+b)f_{n,i-1}.$ (12)
In the following discussion, assume that $a+b\geqslant c\geqslant b>0$. Note that
$f_{1}(x)=c+bx,~f_{2}(x)=c^{2}+(ac+2bc+ab)x+b^{2}x^{2},$
$f_{3}(x)=c^{3}+(a^{2}c+a^{2}b+3abc+3bc^{2}+3ac^{2})x+(a^{2}c+a^{2}b+3abc+3b^{2}c+3ab^{2})x^{2}+b^{3}x^{3}.$
The result is true for $n=1,2,3$, since $b\leqslant c,~b^{2}\leqslant c^{2},~c^{2}\leqslant ac+2bc+ab$ and
$\frac{b^{3}}{c^{3}}\leqslant\frac{a^{2}c+a^{2}b+3abc+3b^{2}c+3ab^{2}}{a^{2}c+a^{2}b+3abc+3bc^{2}+3ac^{2}}\leqslant 1,~\frac{c^{3}}{a^{2}c+a^{2}b+3abc+3b^{2}c+3ab^{2}}\leqslant 1,$
where the last inequality can be derived by using the fact that $(a+b)^{2}\geqslant c^{2}$.
Suppose that $f_{n}(x)$ is the ratio monotone. When $n=2m$, we have
$\frac{f_{2m,2m}}{f_{2m,0}}\leqslant\frac{f_{2m,2m-1}}{f_{2m,1}}\leqslant\cdots\leqslant\frac{f_{2m,2m-i}}{f_{2m,i}}\leqslant\cdots\leqslant\frac{f_{2m,m+1}}{f_{2m,m-1}}\leqslant 1$ (13)
and
$\frac{f_{2m,0}}{f_{2m,2m-1}}\leqslant\frac{f_{2m,1}}{f_{2m,2m-2}}\leqslant\cdots\leqslant\frac{f_{2m,i-1}}{f_{2m,2m-i}}\leqslant\cdots\leqslant\frac{f_{2m,m-1}}{f_{2m,m}}\leqslant 1.$ (14)
We proceed by induction. In the following, we need to show that
$\frac{f_{2m+1,2m+1}}{f_{2m+1,0}}\leqslant\frac{f_{2m+1,2m}}{f_{2m+1,1}}\leqslant\cdots\leqslant\frac{f_{2m+1,2m+1-i}}{f_{2m+1,i}}\leqslant\cdots\leqslant\frac{f_{2m+1,m+1}}{f_{2m+1,m}}\leqslant 1$ (15)
and
$\frac{f_{2m+1,0}}{f_{2m+1,2m}}\leqslant\frac{f_{2m+1,1}}{f_{2m+1,2m-1}}\leqslant\cdots\leqslant\frac{f_{2m+1,i}}{f_{2m+1,2m-i}}\leqslant\cdots\leqslant\frac{f_{2m+1,m-1}}{f_{2m+1,m+1}}\leqslant 1.$ (16)
We first establish (15). Note that
$\Delta_{1}:=f_{2m+1,2m+1}f_{2m+1,1}-f_{2m+1,0}f_{2m+1,2m}$
$=bf_{2m,2m}((a+c)f_{2m,1}+(2ma+b)f_{2m,0})-cf_{2m,0}((2ma+c)f_{2m,2m}+(a+b)f_{2m,2m-1})$
$=$ $(b-c)(2am+b+c)f_{2m,0}f_{2m,2m}+b(a+c)f_{2m,1}f_{2m,2m}-c(a+b)f_{2m,0}f_{2m,2m-1}.$
From the left side of (13), we see that $f_{2m,1}f_{2m,2m}\leqslant f_{2m,0}f_{2m,2m-1}$. It follows from $b\leqslant c$ that
$\Delta_{1}\leqslant(b-c)(2am+b+c)f_{2m,0}f_{2m,2m}+b(a+c)f_{2m,0}f_{2m,2m-1}-c(a+b)f_{2m,0}f_{2m,2m-1}$
$=(b-c)(2am+b+c)f_{2m,0}f_{2m,2m}+a(b-c)f_{2m,2m-1}\leqslant 0.$
Hence
$\frac{f_{2m+1,2m+1}}{f_{2m+1,0}}\leqslant\frac{f_{2m+1,2m}}{f_{2m+1,1}}.$
From the right sides of (13) and (14), we see that $f_{2m,m+1}\leqslant f_{2m,m-1}\leqslant f_{2m,m}$. So we have
$\Delta_{2}:=f_{2m+1,m+1}-f_{2m+1,m}$
$=(a(m+1)+c)f_{2m,m+1}+(ma+b)f_{2m,m}-(am+c)f_{2m,m}-(a(m+1)+b)f_{2m,m-1}$
$=(b-c)f_{2m,m}+(a(m+1)+c)f_{2m,m+1}-(a(m+1)+b)f_{2m,m-1}$
$\leqslant(b-c)f_{2m,m}+(a(m+1)+c)f_{2m,m-1}-(a(m+1)+b)f_{2m,m-1}$
$=(b-c)(f_{2m,m}-f_{2m,m-1})\leqslant 0,$
which yields that $f_{2m+1,m+1}\leqslant f_{2m+1,m}$. We now ready to show that for $1\leqslant i\leqslant m-1$, we have
$\frac{f_{2m+1,2m+1-i}}{f_{2m+1,i}}\leqslant\frac{f_{2m+1,2m-i}}{f_{2m+1,i+1}}.$
From (13) and (14), we observe that
$\frac{f_{2m,2m-i+1}}{f_{2m,i-1}}\leqslant\frac{f_{2m,2m-i}}{f_{2m,i}}\leqslant\frac{f_{2m,2m-i-1}}{f_{2m,i+1}}\leqslant 1,$
$\frac{f_{2m,i-1}}{f_{2m,2m-i}}\leqslant\frac{f_{2m,i}}{f_{2m,2m-i-1}}\leq 1.$
In Lemma 3, setting $a_{1}=f_{2m,2m-i+1},~a_{2}=f_{2m,i-1},~a_{3}=f_{2m,2m-i},~a_{4}=f_{2m,i},~a_{5}=f_{2m,2m-i-1}$, $a_{6}=f_{2m,i+1}$, $\lambda_{1}=a(2m+1-i)+c,~\lambda_{2}=a(2m+1-i)+b,~\lambda=(2m+1)a+b+c$ and $\mu=a$, it follows from (12) that
$\frac{f_{2m+1,2m+1-i}}{f_{2m+1,i}}=\frac{(a(2m+1-i)+c)f_{2m,2m-i+1}+(ai+b)f_{2m,2m-i}}{(a(2m+1-i)+b)f_{2m,i-1}+(ai+c)f_{2m,i}}$
$\leqslant$ $\frac{(a(2m-i)+c)f_{2m,2m-i}+(a(i+1)+b)f_{2m,2m-i-1}}{(a(2m-i)+b)f_{2m,i}+(a(i+1)+c)f_{2m,i+1}}=\frac{f_{2m+1,2m-i}}{f_{2m+1,i+1}}.$
Hence the proof of (15) is complete.
Next, we proceed to prove (16). From (13) and (14), we see that
$f_{2m,0}f_{2m,2m-2}\leqslant f_{2m,1}f_{2m,2m-1},~f_{2m,2m}\leqslant f_{2m,2m-1}\leqslant f_{2m,2m-2},~f_{2m,0}\leqslant f_{2m,1}.$
Thus we get
$\Delta_{3}:=f_{2m+1,0}f_{2m+1,2m-1}-f_{2m+1,1}f_{2m+1,2m}$
$=(c-a-b)(2am+b+c)f_{2m,0}f_{2m,2m-1}-(2am+c)(a+c)f_{2m,1}f_{2m,2m}$
- $(2am+c)(2am+b)f_{2m,0}f_{2m,2m}+c(2a+b)f_{2m,0}f_{2m,2m-2}-(a+c)(a+b)f_{2m,1}f_{2m,2m-1}$
$\leqslant$ $(c-a-b)(2am+b+c)f_{2m,0}f_{2m,2m}-(2am+c)(a+c)f_{2m,1}f_{2m,2m}$
- $(2am+c)(2am+b)f_{2m,0}f_{2m,2m}+a(c-a-b)f_{2m,0}f_{2m,2m-2}.$
Since $a+b\geqslant c$, we get $\Delta_{3}\leqslant 0$. Therefore, we obtain
$\frac{f_{2m+1,0}}{f_{2m+1,2m}}\leqslant\frac{f_{2m+1,1}}{f_{2m+1,2m-1}}.$
From the right sides of (13) and (14), we have $f_{2m,m-2}\leqslant f_{2m,m+1}\leqslant f_{2m,m-1}\leqslant f_{2m,m}$. So we find that
$\Delta_{4}:=f_{2m+1,m-1}-f_{2m+1,m+1}$
$=(am-a+c)f_{2m,m-1}+(am+2a+b)f_{2m,m-2}-(am+a+c)f_{2m,m+1}-(am+b)f_{2m,m}$
$\leqslant(am-a+c)f_{2m,m}+(am+2a+b)f_{2m,m+1}-(am+a+c)f_{2m,m+1}-(am+b)f_{2m,m}$
$=(c-a-b)(f_{2m,m}-f_{2m,m+1})\leqslant 0,$
which yields that $f_{2m+1,m-1}\leqslant f_{2m+1,m+1}$. For $1\leqslant i\leqslant m-2$, we now show that
$\frac{f_{2m+1,i}}{f_{2m+1,2m-i}}\leqslant\frac{f_{2m+1,i+1}}{f_{2m+1,2m-i-1}}.$
From (13) and (14), we have
$\frac{f_{2m,i-1}}{f_{2m,2m-i}}\leqslant\frac{f_{2m,i}}{f_{2m,2m-i-1}}\leqslant\frac{f_{2m,i+1}}{f_{2m,2m-i-2}},$
$\frac{f_{2m,2m-i}}{f_{2m,i}}\leqslant\frac{f_{2m,2m-i-1}}{f_{2m,i+1}}.$
In Lemma 3, setting $a_{1}=f_{2m,i-1},~a_{2}=f_{2m,2m-i},~a_{3}=f_{2m,i}$, $a_{4}=f_{2m,2m-i-1}$, $a_{5}=f_{2m,i+1}$, $a_{6}=f_{2m,2m-i-2}$, $\lambda_{1}=a(2m-i)+a+b$, $\lambda_{2}=a(2m-i)+c$, $\lambda=(2m+1)a+b+c$ and $\mu=a$, it follows from (12) that
$\frac{f_{2m+1,i}}{f_{2m+1,2m-i}}=\frac{(2ma-ai+a+b)f_{2m,i-1}+(ai+c)f_{2m,i}}{(2ma-ai+c)f_{2m,2m-i}+(ai+a+b)f_{2m,2m-i-1}}$
$\leqslant\frac{(2ma-ai+b)f_{2m,i}+(ai+a+c)f_{2m,i+1}}{(2ma-ai-a+c)f_{2m,2m-i-1}+(ai+2a+b)f_{2m,2m-i-2}}=\frac{f_{2m+1,i+1}}{f_{2m+1,2m-i-1}},$
as desired. This completes the proof of (16). The case that $n=2m+1$ can be dealt with in the same manner, and we omit it for simplicity. ∎
3 Applications of Theorem 1
In this section, we apply Theorem 1 to derive certain new results in a unified manner.
3.1 $q$-Eulerian polynomials, $1/k$-Eulerian polynomials and generalized Eulerian polynomials
Let $\mathcal{S}_{n}$ be the set of all permutations of $n]=\{1,2,\ldots,n\}$. For $\pi\in\mathcal{S}_{n}$, we say that $i$ is an excedance if $\pi(i)>i$. Let ${\rm exc\,}(\pi)$ and ${\rm cyc\,}(\pi)$ be the numbers of excedances and cycles of $\pi$, respectively. In [8], Brenti studied the following $q$-Eulerian polynomials:
$A_{n}(x,q)=\sum_{\pi\in\mathcal{S}_{n}}x^{{\rm exc\,}(\pi)}q^{{\rm cyc\,}(\pi)}.$
In particular, $A_{1}(x,q)=q,~A_{2}(x,q)=q(q+x)$ and $A_{3}(x,q)=q(q^{2}+(3q+1)x+x^{2})$. According to [8, Propositions 7.2, 7.3], the $q$-Eulerian polynomials $A_{n}(x,q)$ satisfy the recursion
$A_{n+2}(x,q)=(nx+x+q)A_{n+1}(x,q)+x(1-x)\frac{\partial}{\partial x}A_{n+1}(x,q),$ (17)
and the exponential generating function of these polynomials is given as follows:
$1+\sum_{n\geqslant 1}A_{n}(x,q)\frac{z^{n}}{n!}={\left(\frac{1-x}{\mathrm{e}^{z(x-1)}-x}\right)}^{q}.$
When $q$ is a positive rational number, Brenti showed that $A_{n}(x,q)$ has only real nonpositive simple zeros, and so it is log-concave and unimodal ([8, Theorem 7.5]).
Setting $L_{n}(x)=x^{n}A_{n+1}(1/x,q)$, it follows from (5) and (17) that
$L_{n+1}(x)=(nx+qx+1)L_{n}(x)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}L_{n}(x).$ (18)
Following Hwang-Chern-Duh [17, p. 26], the polynomials $L_{n}(x)$ can be called LI Shanlan polynomials, since these polynomials appeared in his 1867 book. Combining the recursions (17), (18) and Theorem 1, we can give the following result.
Corollary 4.
For any $n\geqslant 1$, we have the following results:
  • ($c_{1}$)
    When $0<q\leqslant 1$, the polynomial $A_{n}(x,q)$ is bi-gamma-positive;
  • ($c_{2}$)
    When $0<q\leqslant 1$, the polynomial $x^{n-1}A_{n}(1/x,q)$ is ratio monotone;
  • ($c_{3}$)
    When $1\leqslant q\leqslant 2$, the polynomial $A_{n}(x,q)$ is ratio monotone and $A_{n}(x,q)$ can be written as a sum of two gamma-positive polynomials with their degrees differing by 1.
It should be noted that $A_{n}(x,2)$ is the big descent polynomials over $\mathfrak{S}_{n+1}$, where a big descent is an index $i$ such that $\pi(i)\geqslant\pi(i+1)+2$, see [27, A120434]. We list the first few polynomials:
$A_{2}(x,2)=4+2x,~A_{3}(x,2)=8+14x+2x^{2},~A_{4}(x,2)=16+66x+36x^{2}+2x^{3}.$
When $n\geqslant 4$, by (6) and (18), it is routine to verify that if $q>2$, then $A_{n}(x,q)$ can not be written as a sum of two gamma-positive polynomials with their degrees differing by 1. For examples,
$A_{4}(x,3)=81+201x+75x^{2}+3x^{3},~A_{4}(x,4)=256+452x+128x^{2}+4x^{3}.$
Following Savage-Viswanathan [24], the $1/k$-Eulerian polynomials $A_{n}^{(k)}(x)$ are defined by
$\sum_{n=0}^{\infty}A_{n}^{(k)}(x)\frac{z^{n}}{n!}=\left(\frac{1-x}{\mathrm{e}^{kz(x-1)}-x}\right)^{\frac{1}{k}},$
where $k\geqslant 1$. They found that $A_{n}^{(k)}(x)$ are the ascent polynomials over $\rm k$-inversion sequences. A more well known interpretation is given as follows (see [24, 25]):
$A_{n}^{(k)}(x)=\sum_{\pi\in\mathcal{S}_{n}}x^{{\rm exc\,}(\pi)}k^{n-{\rm cyc\,}(\pi)}.$ (19)
The polynomials $A_{n}^{(k)}(x)$ are also the ascent-plateau polynomials over $k$-Stirling permutations [20, 22]. They satisfy the recursion
$A_{n+2}^{(k)}(x)=(nkx+kx+1)A_{n+1}^{(k)}(x)+kx(1-x)\frac{\mathrm{d}}{\mathrm{d}x}A_{n+1}^{(k)}(x).$
Below are these polynomials for $n\leqslant 3$:
$A_{1}^{(k)}(x)=1,~A_{2}^{(k)}(x)=1+kx,~A_{3}^{(k)}(x)=1+3kx+k^{2}x(1+x).$
The bi-gamma-positivity of $A_{n}^{(k)}(x)$ was first established in [21, Section 3.4], and Yan-Yang-Lin [30] gave a nice combinatorial proof of this result. Combining Corollary 4 and (19), we get the following result.
Theorem 5.
The reciprocal $1/k$-Eulerian polynomials $x^{n}A_{n+1}^{(k)}(1/x)$ are ratio monotone.
For example, $A_{5}^{(k)}(x)=1+(10k+10k^{2}+5k^{3}+k^{4})x+(25k^{2}+30k^{3}+11k^{4})x^{2}+(15k^{3}+11k^{4})x^{3}+k^{4}x^{4}$. When $k\geqslant 1$, we have
$\frac{1}{k^{4}}\leqslant\frac{10k+10k^{2}+5k^{3}+k^{4}}{15k^{3}+11k^{4}}\leqslant 1,$
$\frac{k^{4}}{10k+10k^{2}+5k^{3}+k^{4}}\leqslant\frac{15k^{3}+11k^{4}}{25k^{2}+30k^{3}+11k^{4}}\leqslant 1.$
Consider a kind of generalized Eulerian polynomials defined by
$P_{n+1}(x;p,q)=(nx+px+q)P_{n}(x;p,q)+x(1-x)\frac{\partial}{\partial x}P_{n}(x;p,q),~P_{0}(x;p,q)=1.$ (20)
These polynomials were introduced by Morisita [23], and they were also studied by Carlitz-Scoville [9]. They also appeared in the context of random staircase tableaux [16]. Hitczenko-Janson [16] investigated the asymptotic distribution of these polynomials. Combining (20) and Theorem 1, we end this subsection by giving the following result.
Corollary 6.
If $1+q\geqslant p\geqslant q>0$, then $P_{n}(x;p,q)$ is bi-gamma-positive and $x^{n}P_{n}(1/x;p,q)$ is ratio monotone. If $1+p\geqslant q\geqslant p>0$, then $P_{n}(x;p,q)$ is ratio monotone.
3.2 The $q$-Eulerian polynomials of type $B$
Let ${\mathcal{S}}^{B}_{n}$ denote the hyperoctahedral group of rank $n$. Elements of ${\mathcal{S}}^{B}_{n}$ are signed permutations of the set $\pm[n]=[n]\cup\{\overline{1},\ldots,\overline{n}\}$ with the property that $\sigma(\overline{i})=-\sigma(i)$ for all $i\in[n]$, where $\overline{i}=-i$. Following Brenti [7], the $q$-Eulerian polynomials of type $B$ are defined by
$B_{n}(x,q)=\sum_{\sigma\in{\mathcal{S}}^{B}_{n}}x^{\operatorname{des}_{B}(\sigma)}q^{\operatorname{neg}(\sigma)},$
where $\operatorname{neg}(\sigma)=\#\{i\in[n]:~\pi(i)<0\}$ and
$\operatorname{des}_{B}(\pi)=\#\{i\in\{0,1,2,\ldots,n-1\}:~\pi(i)>\pi({i+1})~\&~\pi(0)=0\}.$
They satisfy the recursion
$B_{n+1}(x,q)=((1+q)nx+qx+1)B_{n}(x,q)+(1+q)x(1-x)\frac{\partial}{\partial x}B_{n}(x,q),$
with the initial conditions $B_{0}(x,q)=1$ and $B_{1}(x,q)=1+qx$. Setting $h_{n}(x)=x^{n}B_{n}(1/x,q)$, it follows from (5) that
$h_{n+1}(x)=((1+q)nx+x+q)h_{n}(x,q)+(1+q)x(1-x)\frac{\partial}{\partial x}h_{n}(x,q).$
Combining the above the recursions and Theorem 1, we get the following result.
Corollary 7.
For any $n\geqslant 1$, we have the following results:
  • ($c_{1}$)
    When $q\geqslant 1$, the polynomial $B_{n}(x,q)$ is bi-gamma-positive;
  • ($c_{2}$)
    When $q\geqslant 1$, the polynomial $x^{n}B_{n}(1/x,q)$ is ratio monotone;
  • ($c_{3}$)
    When $0<q\leqslant 1$, the polynomial $B_{n}(x,q)$ is ratio monotone and it can be written as a sum of two gamma-positive polynomials with their degrees differing by 1.
For example, $B_{4}(x,q)=1+(11+32q+24q^{2}+8q^{3}+q^{4})x+(11+56q+96q^{2}+56q^{3}+11q^{4})x^{2}+(1+8q+24q^{2}+32q^{3}+11q^{4})x^{3}+q^{4}x^{4}$. When $q\geqslant 1$, we have
$\frac{1}{q^{4}}\leqslant\frac{11+32q+24q^{2}+8q^{3}+q^{4}}{1+8q+24q^{2}+32q^{3}+11q^{4}}\leqslant 1;$
$\frac{q^{4}}{11+32q+24q^{2}+8q^{3}+q^{4}}\leqslant\frac{1+8q+24q^{2}+32q^{3}+11q^{4}}{11+56q+96q^{2}+56q^{3}+11q^{4}}\leqslant 1.$
3.3 The $r$-colored Eulerian polynomials
Following Steingrímsson [28], the $r$-colored Eulerian polynomial can be defined by
$A_{n+1,r}(x)=(rnx+(r-1)x+1)A_{n,r}(x)+rx(1-x)\frac{\mathrm{d}}{\mathrm{d}x}A_{n,r}(x),~A_{0,r}(x)=1.$ (21)
When $r=1$ and $r=2$, the polynomial $A_{n,r}(x)$ reduces to the types $A$ and $B$ Eulerian polynomials $A_{n}(x)$ and $B_{n}(x)$, respectively. Very recently, there has been much work devoted to the bi-gamma-positivity of the $r$-colored Eulerian polynomials and their variations, see [2, 15, 21] for details. In particular, it is now well known that $A_{n,r}(x)$ is bi-gamma-positive when $r>2$, see [2, Eq. (21)] and [21, Theorem 7.5].
Setting $e_{n,r}(x)=x^{n}A_{n,r}(1/x)$, it follows from (5) and (3.3) that
$e_{n+1,r}(x)=(rnx+x+r-1)e_{n,r}(x)+rx(1-x)\frac{\mathrm{d}}{\mathrm{d}x}e_{n,r}(x),~e_{0,r}(x)=1.$
Therefore, by Theorem 1, we arrive at the following result.
Corollary 8.
For any $n\geqslant 1$, we have the following results:
  • ($c_{1}$)
    When $r\geqslant 2$, the reciprocal polynomial $x^{n}A_{n,r}(1/x)$ is ratio monotone;
  • ($c_{2}$)
    When $1\leqslant r\leqslant 2$, the polynomial $A_{n,r}(x)$ is ration monotone, and it can be written as a sum of two gamma-positive polynomials with their degrees differing by 1.
Acknowledgements
The second author is supported by the National Natural Science Foundation of China (No. 12401459) and the Natural Science Foundation of Shandong Province of China (ZR2024QA075). The third author is supported by the National Natural Science Foundation of China (No. 12071063) and Taishan Scholars Program of Shandong Province (No. tsqn202211146).

### v3

@@
 MSC:
-[2010] 05A05, 05A15
+[2010] 26D05, 05A15
 1 Introduction
@@
 Clearly, a log-concave sequence with no internal zeros is unimodal, but the converse is not true.
-The polynomial $a(x)$ is said to be gamma-positive if there exist nonnegative numbers $\gamma_{k}$ such that
+Darroch [14] showed that if $a(x)$ has only real nonpositive zeros, then $a(x)$ is unimodal with at most two modes and each mode $k$ satisfies
+$\left\lfloor\frac{a^{\prime}(1)}{a(1)}\right\rfloor\leqslant k\leqslant\left\lceil\frac{a^{\prime}(1)}{a(1)}\right\rceil.$
+The polynomial $a(x)$ is called gamma-positive if there exist nonnegative numbers $\gamma_{k}$ such that
 $a(x)=\sum_{k=0}^{\lfloor{n}/{2}\rfloor}\gamma_{k}x^{k}(1+x)^{n-2k}.$
@@
 We say that $a(x)$ is spiral if $a_{n}\leqslant a_{0}\leqslant a_{n-1}\leqslant a_{1}\leqslant\cdots\leqslant a_{\lfloor n/2\rfloor}$.
-Following [26, Definition 2.9], the polynomial $a(x)$ is alternatingly increasing if $a_{0}\leqslant a_{n}\leqslant a_{1}\leqslant a_{n-1}\leqslant\cdots\leqslant a_{\lfloor{(n+1)}/{2}\rfloor}$.
+Following [27, Definition 2.9], the polynomial $a(x)$ is alternatingly increasing if $a_{0}\leqslant a_{n}\leqslant a_{1}\leqslant a_{n-1}\leqslant\cdots\leqslant a_{\lfloor{(n+1)}/{2}\rfloor}$.
 Clearly, if $a(x)$ is spiral and $\deg a(x)=n$, then $x^{n}a(1/x)$ is alternatingly increasing, and vice versa.
@@
 Following Brändén-Solus [5], we call the ordered pair of polynomials $(a(x),b(x))$ the symmetric decomposition of $f(x)$.
-According to [21, Definition 1.2], if $a(x)$ and $b(x)$ are both gamma-positive, then $f(x)$ is said to be bi-gamma-positive.
+According to [22, Definition 1.2], if $a(x)$ and $b(x)$ are both gamma-positive, then $f(x)$ is said to be bi-gamma-positive.
 Bi-gamma-positivity implies alternatingly increasing property and gamma-positivity can be seen as a degeneration of bi-gamma-positivity.
@@
 If $a+c\geqslant b\geqslant c>0$, then $f_{n}(x)$ is bi-gamma-positive and $x^{n}f_{n}(1/x)$ is ratio monotone.
+As Hwang-Chern-Duh [18, Section 4.1] once said, one of the most common recursions with rich combinatorial properties among the extensions of Eulerian numbers is given as follows:
+$\mathcal{P}_{n+1}(x)=(qnx+(qr-p)x+p)\mathcal{P}_{n}(x)+qx(1-x)\frac{\mathrm{d}}{\mathrm{d}x}\mathcal{P}_{n}(x),~\mathcal{P}_{0}(x)=1,$ (5)
+which covers more then 60 examples in OEIS (see [18, Section 4]). The exponential generating function of $\mathcal{P}_{n}(x)$ has the closed-form [18, Eq. (36)]:
+$\sum_{n=0}^{\infty}\mathcal{P}_{n}(x)\frac{z^{n}}{n!}=\mathrm{e}^{p(1-x)z}\left(\frac{1-x}{1-x\mathrm{e}^{q(1-x)z}}\right)^{r}.$
+Using Theorem (1) and the recursions (5) and (6), we get the following result.
+Corollary 2.
+Let $\mathcal{P}_{n}(x)$ be the polynomials satisfying the recursion (5).
+If $q+2p\geqslant qr\geqslant 2p>0$, then $\mathcal{P}_{n}(x)$ is bi-gamma-positive and $x^{n}\mathcal{P}_{n}(1/x)$ is ratio monotone.
+If $q(1+r)\geqslant 2p\geqslant qr>0$, then $\mathcal{P}_{n}(x)$ is ratio monotone.
+This paper is organized as follows. Section 2 is devoted to the proof of Theorem 1. In Section 3, we present some applications of Theorem 1.
 2 Proof of Theorem 1
@@
 with $\alpha_{0,0}=1,~\beta_{0,0}=0$ and $\alpha_{0,k}=\beta_{0,k}=0$ for $k\geqslant 1$.
-Therefore, when $a+c\geqslant b\geqslant c>0$, we see that $f_{n}(x)$ is bi-gamma-positive, and when $a>0$ and $b=c$, $f_{n}(x)$ is gamma-positive.
+Therefore, when $a+c\geqslant b\geqslant c>0$, we see that $f_{n}(x)$ is bi-gamma-positive.
+Moreover, if $a>0$ and $b=c$, then $f_{n}(x)$ is reduced to a gamma-positive polynomial.
 ∎
@@
 We use the notation $p(x)\prec q(x)$ for $p(x)$ interlaces $q(x)$.
-Combining (6) and [18, Theorem 2.3], one can immediately get the following result.
+By [19, Theorem 2.1], one can immediately get the following result.
 Proposition 3.
-Let $\{f_{n}(x)\}_{n\geqslant 0}$ be a sequence of polynomials satisfying (4).
-Then $f_{n}(x)$ has only real zeros.
-Suppose that $(a_{n}(x),b_{n}(x))$ is the symmetric decomposition of $f_{n}(x)$.
-Then both $a_{n}(x)$ and $b_{n}(x)$ have only real zeros and $b_{n}(x)\prec a_{n}(x)$.
+Let $\{f_{n}(x)\}_{n\geqslant 0}$ be a sequence of polynomials satisfying (4). Then $f_{n}(x)$ has only real zeros and so it is unimodal.
 2.2 A proof of the ratio monotonicity of $f_{n}(x)$
@@
 $\frac{\lambda_{1}a_{1}+(\lambda-\lambda_{1})a_{3}}{\lambda_{2}a_{2}+(\lambda-\lambda_{2})a_{4}}\leqslant\frac{\lambda_{1}a_{3}+(\lambda-\lambda_{1})a_{5}}{\lambda_{2}a_{4}+(\lambda-\lambda_{2})a_{6}}\leqslant\frac{a_{5}-a_{3}}{a_{6}-a_{4}}.$
-Comparing with (8) and (9), we immediately get the desired result. This completes the proof. ∎
+Comparing with (9) and (10), we get the desired result. This completes the proof. ∎
 A proof of the ratio monotonicity of $f_{n}(x)$:
@@
 When $q$ is a positive rational number, Brenti showed that $A_{n}(x,q)$ has only real nonpositive simple zeros, and so it is log-concave and unimodal ([8, Theorem 7.5]).
-Setting $L_{n}(x)=x^{n}A_{n+1}(1/x,q)$, it follows from (5) and (17) that
-$L_{n+1}(x)=(nx+qx+1)L_{n}(x)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}L_{n}(x).$ (18)
-Following Hwang-Chern-Duh [17, p.
-26], the polynomials $L_{n}(x)$ can be called LI Shanlan polynomials, since these polynomials appeared in his 1867 book.
+Setting $L_{n}(x,q)=x^{n}A_{n+1}(1/x,q)$, it follows from (6) and (18) that
+$L_{n+1}(x,q)=(nx+qx+1)L_{n}(x,q)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}L_{n}(x,q).$ (19)
+Following Hwang-Chern-Duh [18, p.
+26], the polynomials $L_{n}(x,q)$ can be called LI Shanlan polynomials, since these polynomials first appeared in his 1867 book.
 Combining the recursions (18), (19) and Theorem 1, we can give the following result.
@@
     When $1\leqslant q\leqslant 2$, the polynomial $A_{n}(x,q)$ is ratio monotone and $A_{n}(x,q)$ can be written as a sum of two gamma-positive polynomials with their degrees differing by 1.
-It should be noted that $A_{n}(x,2)$ is the big descent polynomials over $\mathfrak{S}_{n+1}$, where a big descent is an index $i$ such that $\pi(i)\geqslant\pi(i+1)+2$, see [27, A120434].
+It should be noted that $A_{n}(x,2)$ is the big descent polynomials over $\mathfrak{S}_{n+1}$, where a big descent is an index $i$ such that $\pi(i)\geqslant\pi(i+1)+2$, see [28, A120434].
 We list the first few polynomials:
@@
 $A_{1}^{(k)}(x)=1,~A_{2}^{(k)}(x)=1+kx,~A_{3}^{(k)}(x)=1+3kx+k^{2}x(1+x).$
-The bi-gamma-positivity of $A_{n}^{(k)}(x)$ was first established in [21, Section 3.4], and Yan-Yang-Lin [30] gave a nice combinatorial proof of this result.
+The bi-gamma-positivity of $A_{n}^{(k)}(x)$ was first established in [22, Section 3.4], and Yan-Yang-Lin [31] gave a nice combinatorial proof of this result.
 Combining Corollary 5 and (20), we get the following result.
@@
 $P_{n+1}(x;p,q)=(nx+px+q)P_{n}(x;p,q)+x(1-x)\frac{\partial}{\partial x}P_{n}(x;p,q),~P_{0}(x;p,q)=1.$ (21)
-These polynomials were introduced by Morisita [23], and they were also studied by Carlitz-Scoville [9].
-They also appeared in the context of random staircase tableaux [16].
-Hitczenko-Janson [16] investigated the asymptotic distribution of these polynomials.
+These polynomials were introduced by Morisita [24], and they were also independently studied by Carlitz-Scoville [9].
+It should be noted that these polynomials appeared in the context of random staircase tableaux, and Hitczenko-Janson [17] investigated their asymptotic distribution.
 Combining (21) and Theorem 1, we end this subsection by giving the following result.
@@
 In particular, it is now well known that $A_{n,r}(x)$ is bi-gamma-positive when $r>2$, see [2, Eq.
-(21)] and [21, Theorem 7.5].
+(21)] and [22, Theorem 7.5].
 Setting $e_{n,r}(x)=x^{n}A_{n,r}(1/x)$, it follows from (6) and (3.3) that